OpenStudy (lgbasallote):

Please help (Integral Calculus question) Find the area bounded by the curve \(y = 5x - x^2\) and the line \(y = x\) How do I do this??

OpenStudy (anonymous):

k I'm taking this as I go so give me a second first you need to find the points of intersection between the two lines do you know how to do this?

OpenStudy (lgbasallote):

no...how?

OpenStudy (lgbasallote):

well it is given in my book...but i'd like to know as well how to find it without its being given

OpenStudy (anonymous):

don't you just set the equations equal to one another and then solve? because basically you want to know what value of x gives the same value of y for both equations right?

OpenStudy (anonymous):

well what values I should say

OpenStudy (lgbasallote):

ahh so substitution method...why didn't i think of that \[x = 5x - x^2\] \[0 = 4x - x^2\] \[0 = x(4 - x)\] x =0 x = 4 these are the intersections right? so point (0,0) and (4,4)....exactly what it says in my book thanks

OpenStudy (lgbasallote):

so what's next after the points of intersection?

OpenStudy (anonymous):

okay you need to use a formula there aren't different integration formulas for different situations but this one was straight forward luckily. First for the interval x= 0 to 4 what function is bigger?

OpenStudy (anonymous):

there are different** sorry typo

OpenStudy (lgbasallote):

well seems y = 5x - x^2

OpenStudy (anonymous):

did you check? I haven't done the math but you should probably check do you know how to?

OpenStudy (anonymous):

It is really easy you just have to put any value of x in btw 0 and 4 into both equations and see which is bigger I used 1 and it turns out you were right haha it was pretty obvious but when both the equations are exponential it won't be so obvious.

OpenStudy (lgbasallote):

 that's the graph